Part A What is the speed of an electron with a de Broglie wavelength of 0.17nm...

Part A

What is the speed of an electron with a de Broglie wavelength of 0.17nm ?

Express your answer using two significant figures.

Part B

What is the speed of a proton with a de Broglie wavelength of 0.17nm ?

Express your answer using two significant figures.

Part A What is the speed of an electron with a de
Answer #1

De Broglie wavelength is given by

P= \frac{h}{\lambda }

mV= \frac{h}{\lambda }

V= \frac{h}{m\times \lambda }

(A) for electron

h is planck constant h= 6.626